OpenStudy (anonymous):

Solve for x: 3x − 24 = 81 A. fifty-seven over three B. one hundred five over three C. 57 D. 105 PLEASE HELP

OpenStudy (anonymous):

@superman36

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

first add \(24\)

OpenStudy (misty1212):

then divide by 3

OpenStudy (spring98):

https://mathway.com/

OpenStudy (anonymous):

what is the answer

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

@misty1212

OpenStudy (anonymous):

please help

OpenStudy (misty1212):

what is \(81+24\)?

OpenStudy (anonymous):

105

OpenStudy (misty1212):

right so \[3x-21=81\\ 3x=105\] now divide by \(3\)

OpenStudy (misty1212):

just leave it as a fraction \[3x=105\\ x=\frac{105}{3}\] which looks like choice 2

OpenStudy (anonymous):

what ?

OpenStudy (misty1212):

isn't choice two \[\frac{105}{3}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is that it

OpenStudy (misty1212):

yes halas